Farrell wants to retire in six years. To have sufficient assets to fund retirement, Farrell needs to accumulate an additional $400,000 between today and retirement. As his planner, you assume that inflation will average 5%. You are also confident that you can build a portfolio that will generate an 8% compounded annual after-tax return. What serial payment should Farrell invest at the end of the first year to fund this goal?
Answer #1
Amount needed at the end of 6 years after considering inflation = 400000*1.05^6 = $ 5,36,038
This $536,038 is the FV of the serial deposits
(Annuity) to be made at the end of each year for
the next 6 years.
Hence, 536038 = A*(1.08^6-1)/0.08 (using the
formula for FV of an annuity), where A is the
annual investment to be made.
Annual investment from the end of the first year = 536038*0.08/(1.08^6-1) = $     73,070